a particle starting from rest and moving with an acceleration 4/3m/s-2 , what is the distance travelled by particle in nth second?

Grade:12th pass

1 Answers

Soumya Ranjan Mohanty
6 years ago
Distance travelled in nth second=Sn=u+a/2(2n-1) Here a = 4/3 m/s^2 and u=0.
so Sn=0+ (4/3)/2 * (2n-1) =2/3 (2n-1) unit distance.